# Fractional Sobolev paths on Wasserstein spaces and their energy-minimizing particle representations

### v3

@@ 2.7. Piecewise geodesic approximation on geodesic spaces
$D_{n}\coloneqq\{0=t_{0}^{(n)}<t_{1}^{(n)}<\cdots<t_{N_{n}}^{(n)}=1\},$ (2.53)
$t_{i}^{(n)}\coloneqq\frac{i}{2^{n}}\quad i\in\left\{0,1,\cdots,2^{n}\eqqcolon N_{n}\right\}.$ (2.54)
@@
$|X^{n}|_{b^{\alpha,p}}^{p}=\sum_{m=0}^{n}2^{m(\alpha p-1)}\sum_{\begin{subarray}{c}i=k\frac{2^{n}}{2^{m}}\\ k\in\{0,1,\cdots,2^{m}-1\}\end{subarray}}d(x_{i},x_{i+\frac{2^{n}}{2^{m}}})^{p}+\frac{2^{n(\alpha p-1)}}{2^{(p-\alpha p)}-1}\sum_{i=0}^{2^{n}-1}d(x_{i},x_{i+1})^{p}.$ (2.57)
@@
$=\sum_{m=0}^{n}2^{m(\alpha p-1)}\sum_{\begin{subarray}{c}i=k\frac{2^{n}}{2^{m}}\\ k\in\{0,1,\cdots,2^{m}-1\}\end{subarray}}d(x_{i},x_{i+\frac{2^{n}}{2^{m}}})^{p}.$ (2.60)
@@ 2.8. Path measures and sample path regularity
$\int_{\mathcal{X}^{I}}|\gamma|_{\upgamma\textrm{-}\mathrm{H\ddot{o}l}}{\mathrm{\,d}}\pi(\gamma)\overset{?}{<}+\infty$
[…]
@@
In contrast to the functions above, the function $\gamma\mapsto|\gamma|_{b^{\alpha,p}}$ has the advantage of measurability. Recall that its definition only relies on a countable sum of distances. As a consequence of Beppo Levi’s lemma, we can affirm that $\gamma\mapsto|\gamma|^{p}_{b^{\alpha}}$ is measurable because it is a pointwise limit of a non-decreasing sequence of measurable functions:
@@ 2.9. Probability measures, narrow convergence, and tightness
Narrow convergence can be characterized by a subset of $C_{b}(\mathcal{X})$, namely, the space of Lipschitz bounded functions $\mathrm{Lip}_{b}(\mathcal{X})$ (see e.g. [2, Lemma 8.12]). We also recall that if $(\mu_{n})\subset P(\mathcal{X})$ narrowly converges to $\mu\in P(\mathcal{X})$ and $\varphi:\mathcal{X}\to[0,+\infty]$ is a lower semi-continuous function, then
@@ 2.10. Tightness conditions for path measures
Let’s observe that a suitable Hölder regularity can guarantee the finiteness of the second part of (2.76). We thus arrive at the well-known Kolmogorov–Lamperti tightness result (cf. also [10, Corollary A.11]):
@@ 2.12. Compatible measures in $P_{p}(\mathcal{X})$
The compatibility property was implicitly defined in [7, Section 4] as an admissible property for a collection of transport maps and was further explored in [19]. What we outlined above can be generalized to more than 3 measures and to general spaces, for which we give the following definition, mentioned in [19, Remark 2.3.2]. Now, let $(\mathcal{X},d)$ be a complete separable metric space and $p\in[1,\infty)$.
@@
A prominent and well-known example of compatible measures is Wasserstein geodesics, as highlighted in the remark below. A compatible collection, however, need not lie on a Wasserstein geodesic. Take for instance an arbitrary measure and translate it along a non-geodesic curve. Another example is any collection of measures in $P_{p}(\mathbb{R})$ is compatible. As shown in [19, page 49], any collection of Gaussian measures on $\mathbb{R}^{\mathrm{d}}$ that includes the standard Gaussian distribution and such that the covariance matrices of the measures are simultaneously diagonalizable (i.e. one can diagonalize all with one orthogonal matrix) is compatible. See more nontrivial examples presented in [7, Proposition 4.1] and explained in [19, Section 2.3.2].
@@
If $|j-i|=1$ or 0, nothing is needed to prove. Let’s consider $j-i>1$. By triangle inequality and Minkowski’s inequality for $L^{p}$ functions, we obtain
@@ 2.13. Further remarks
with a measurable geodesic selection map $\ell:\mathcal{X}\times\mathcal{X}\to Geo([0,1];\mathcal{X})$. We briefly emphasize that such a map exists and that the above is well-defined (using a similar argument as in the proof of [14, Proposition 1]). Define the multi-valued map $L:\mathcal{X}\times\mathcal{X}\to 2^{Geo([0,1];\mathcal{X})}$ as follows:
@@
As $\mathcal{G}(L)$ is closed, it is Borel-measurable, i.e., it belongs to $\mathcal{B}(\mathcal{X}^{2}\times C)$, which is equal to $\mathcal{B}(\mathcal{X}^{2})\otimes\mathcal{C}$, since all spaces are separable here. Obviously, $\mathcal{G}(L)$ is also in $\mathcal{B}(\mathcal{X}^{2})_{\Upsilon}\otimes\mathcal{C}$, where $\mathcal{B}(\mathcal{X}^{2})_{\Upsilon}$ denotes the $\Upsilon$-completion of $\mathcal{B}(\mathcal{X}^{2})$. It follows from a measurable selection theorem by Aumann (see, e.g., [8, Theorem III.22] and [3, page 455]) that there exists a measurable map $\ell:\big{(}\mathcal{X}^{2},\mathcal{B}(\mathcal{X}^{2})_{\Upsilon}\big{)}\to\big{(}C([0,1];\mathcal{X}),\mathcal{C}\big{)}$ such that $\ell(x,y)\in Geo([0,1];\mathcal{X})$ for $\Upsilon$-a.e. $(x,y)\in\mathcal{X}^{2}$. The above argument for selecting geodesics between two points can easily be generalized to multiple points.
@@ 3.1. Existence of a minimizer for the variational problem
Since the sublevels of $\Psi$ are compact (which follows from lower semi-continuity and relative compactness), the condition (3.2) implies that the family of measures $\{\pi_{n}\}_{n}\subset P(\Gamma_{T})$ is tight by Lemma 2.36. By Prokhorov’s theorem, there exists a subsequence $\{\pi_{n_{k}}\}_{k\in\mathbb{N}}$ such that $\pi_{n_{k}}\to\pi$ narrowly on $P(\Gamma_{T})$ as $k\to\infty$. By Lemma 2.43, we have $\pi\in\mathrm{Lift}(\mu_{t})$ and thus $\int_{\Gamma_{T}}\Psi{\mathrm{\,d}}\pi\geq M$. On the other hand, since $\Psi$ is lower semi-continuous, we obtain the reverse inequality
@@ 3.2. From path measures to Wasserstein curves
Next, we show that $t\mapsto\mu_{t}$ is a continuous curve in $p$-Wasserstein space. Fix an arbitrary $t\in[0,T]$ and take a sequence $t_{n}\to t$ as $n\to\infty$. This sequence will eventually fall in $\delta$-neighbourhood of $t$. Thus, by Lebesgue’s dominated convergence theorem on the measure space $(\Gamma_{T},\mathcal{B}(\Gamma_{T}),\pi)$ and the integrability assumption (3.5), we obtain
@@
where we used Tonelli’s theorem to interchange integrals. Similarly, we have (when $T=1$) that
@@
where we used Beppo Levi’s lemma to interchange limit and integral. ∎
@@
by Lebesgue’s dominated convergence theorem, where the dominated function
@@ 3.3. From Wasserstein curves to path measures: a superposition principle
$(\text{Pr}^{i,i+\frac{2^{n}}{2^{m}}})_{\#}\Upsilon_{n}\in\text{OptCpl}\big{(}\mu_{t^{(n)}_{i}},\mu_{t^{(n)}_{i+\frac{2^{n}}{2^{m}}}}\big{)},$ (3.54)
@@
$=\int_{\bm{\mathcal{X}}_{n}}\Big{(}\sum_{m=0}^{n}2^{m(\alpha p-1)}\sum_{\begin{subarray}{c}i=k\frac{2^{n}}{2^{m}}\\ k\in\{0,1,\cdots,2^{m}-1\}\end{subarray}}d(x_{i},x_{i+\frac{2^{n}}{2^{m}}})^{p}+\frac{2^{n(\alpha p-1)}}{2^{(p-\alpha p)}-1}\sum_{i=0}^{2^{n}-1}d(x_{i},x_{i+1})^{p}\Big{)}{\mathrm{\,d}}\Upsilon_{n}$ (3.61)
$=\sum_{m=0}^{n}2^{m(\alpha p-1)}\sum_{\begin{subarray}{c}i=k\frac{2^{n}}{2^{m}}\\ k\in\{0,1,\cdots,2^{m}-1\}\end{subarray}}\int d(x_{i},x_{i+\frac{2^{n}}{2^{m}}})^{p}{\mathrm{\,d}}\Upsilon_{n}+\frac{2^{n(\alpha p-1)}}{2^{(p-\alpha p)}-1}\sum_{i=0}^{2^{n}-1}\int d(x_{i},x_{i+1})^{p}{\mathrm{\,d}}\Upsilon_{n}$ (3.62)
$=\sum_{m=0}^{n}2^{m(\alpha p-1)}\sum_{\begin{subarray}{c}i=k\frac{2^{n}}{2^{m}}\\ k\in\{0,1,\cdots,2^{m}-1\}\end{subarray}}W_{p}^{p}(\mu_{t^{(n)}_{i}},\mu_{t^{(n)}_{i+2^{n}/2^{m}}})+\frac{2^{n(\alpha p-1)}}{2^{(p-\alpha p)}-1}\sum_{i=0}^{2^{n}-1}W_{p}^{p}(\mu_{t^{(n)}_{i}},\mu_{t^{(n)}_{i+1}})$ (3.63)
@@
Let us clarify the computation steps. After applying the push-forward (3.57) in the first step (3.60), we note that for fixed $n$ and $\bm{x}=(x_{0},\cdots,x_{N_{n}})$, the curve $t\mapsto\ell_{\bm{x}}(t)$ is piecewise geodesic. We have previously computed the $b^{\alpha,p}$-semi-norm of such curves in (2.57), which we now use to obtain (3.61). The compatibility property (3.54) is used to go from (3.62) to (3.63), where even the 2-D marginals of $\Upsilon_{n}$ that are separated by more than one step are given by the Wasserstein distance. The inequality in (3.65) is justified by extending the upper limit of the first sum in (3.64) to infinity, which then corresponds exactly to the definition $|\mu|^{p}_{b^{\alpha,p}}$. The second sum is nothing but one summand in the definition of $|\mu|^{p}_{b^{\alpha,p}}$, so it can be bounded accordingly. To bound (3.66), we use the equivalence (2.23) of the semi-norms $|\cdot|_{W^{\alpha,p}}$ and $|\cdot|_{b^{\alpha,p}}$. The last expression in (3.67) is indeed finite by assumption.
@@
$\int_{\Gamma}\phi(\gamma_{t}){\mathrm{\,d}}\pi(\gamma)\,$ =$\overset{(a)}{=}\lim_{k\to\infty}\int_{\Gamma}\phi(\gamma_{[2^{n_{k}}t]/2^{n_{k}}}){\mathrm{\,d}}\pi_{n_{k}}(\gamma)$ (3.71)
=$\overset{(b)}{=}\lim_{k\to\infty}\int_{\bm{\mathcal{X}}_{n_{k}}}\phi(x_{[2^{n_{k}}t]}){\mathrm{\,d}}\Upsilon_{n_{k}}(\bm{x})$ (3.72)
=$\overset{(c)}{=}\lim_{k\to\infty}\int_{\mathcal{X}}\phi(x){\mathrm{\,d}}\mu_{[2^{n_{k}}t]/2^{n_{k}}}(x)$ (3.73)
=$\overset{(d)}{=}\int_{\mathcal{X}}\phi(x){\mathrm{\,d}}\mu_{t}(x).$ (3.74)
@@
where $d(\gamma_{t},\gamma_{[2^{n_{k}}t]/2^{n_{k}}})$ is estimated using Garsia–Rodemich–Rumsey inequality as described in (2.29) and we used Jensen’s inequality for the integral. Note that thanks to the calculations in the proof of tightness, we now know that $\int|\gamma|^{p}_{W^{\alpha,p}}{\mathrm{\,d}}\pi_{n_{k}}$ is uniformly finite, which means that $\pi_{n_{k}}$-a.e. $\gamma$ has finite $W^{\alpha,p}$-semi-norm. In the limit, the last expression above approaches zero since
@@
Steps $(b)$-$(c)$ simply follow from the construction. More precisely, for fixed $k$, the 1-D marginals of $\pi_{n_{k}}$ at time $t$ coincide with $\mu_{t}$ whenever $t$ is of the form $t=\frac{i}{2^{n_{k}}}$ for some integer $i$.
@@
=$\overset{(c)}{=}\liminf_{m\to\infty}W_{p}^{p}\big{(}\mu_{[2^{m}t]/2^{m}},\mu_{[2^{m}s]/2^{m}}\big{)}$ (3.81)
=$\overset{(d)}{=}W_{p}^{p}(\mu_{t},\mu_{s}),$ (3.82)
where in step $(a)$, we applied Fatou’s lemma on the measure space $(\Gamma,\mathcal{B}(\Gamma),\pi)$. Notice that for every continuous curve $\gamma\in\Gamma$, we have
@@
where we used Tonelli’s theorem. Similarly, we have
@@
where we used Beppo Levi’s lemma. This completes the proof. ∎

### v4

@@ 2.7. Piecewise geodesic approximation on geodesic spaces
$D_{n}\coloneqq\{0=t_{0}^{(n)}<t_{1}^{(n)}<\cdots<t_{N_{n}}^{(n)}=1\},$
$t_{i}^{(n)}\coloneqq\frac{i}{2^{n}}\quad i\in\left\{0,1,\cdots,2^{n}\eqqcolon N_{n}\right\}.$
@@
$|X^{n}|_{b^{\alpha,p}}^{p}=\sum_{m=0}^{n}2^{m(\alpha p-1)}\hskip-20.0pt\sum_{\begin{subarray}{c}i=k\frac{2^{n}}{2^{m}}\\ k\in\{0,1,\cdots,2^{m}-1\}\end{subarray}}\hskip-20.0ptd(x_{i},x_{i+\frac{2^{n}}{2^{m}}})^{p}+\frac{2^{n(\alpha p-1)}}{2^{(p-\alpha p)}-1}\sum_{i=0}^{2^{n}-1}d(x_{i},x_{i+1})^{p}.$ (2.55)
@@
$=\sum_{m=0}^{n}2^{m(\alpha p-1)}\hskip-20.0pt\sum_{\begin{subarray}{c}i=k\frac{2^{n}}{2^{m}}\\ k\in\{0,1,\cdots,2^{m}-1\}\end{subarray}}\hskip-20.0ptd(x_{i},x_{i+\frac{2^{n}}{2^{m}}})^{p}.$ (2.58)
@@ 2.8. Path measures and sample path regularity
$\int_{\mathcal{X}^{I}}|\gamma|_{\upgamma\textrm{-}\mathrm{H\ddot{o}l}}{\mathrm{\,d}}\pi(\gamma)\overset{?}{<}$$\infty$
because $\gamma\mapsto|\gamma|_{\upgamma\textrm{-}\mathrm{H\ddot{o}l}}$ is not measurable as a function from $(\mathcal{X}^{I},\mathcal{B}(\mathcal{X})^{I})\to(\bar{\mathbb{R}}_{+},\mathcal{B}(\bar{\mathbb{R}}_{+}))$. Here $\bar{\mathbb{R}}_{+}\coloneqq[0,+\infty]$ and the corresponding $\sigma$-algebra $\mathcal{B}(\bar{\mathbb{R}}_{+})$ also contains $\{+\infty\}$. Let us summarize this observation:
@@
In contrast to the functions above, the function $\gamma\mapsto|\gamma|_{b^{\alpha,p}}$ has the advantage of measurability. Recall that its definition only relies on a countable sum of distances. As a consequence of Beppo Levi's lemma, we can affirm that $\gamma\mapsto|\gamma|^{p}_{b^{\alpha}}$ is measurable because it is a pointwise limit of a non-decreasing sequence of measurable functions:
@@ 2.9. Probability measures, narrow convergence, and tightness
Narrow convergence can be characterized by a subset of $C_{b}(\mathcal{X})$, namely, the space of Lipschitz bounded functions $\mathrm{Lip}_{b}(\mathcal{X})$ (see e.g. [4, Lemma 8.12]). We also recall that if $(\mu_{n})\subset P(\mathcal{X})$ narrowly converges to $\mu\in P(\mathcal{X})$ and $\varphi:\mathcal{X}\to[0,+\infty]$ is a lower semi-continuous function, then
@@ 2.10. Tightness conditions for path measures
Let's observe that a suitable Hölder regularity can guarantee the finiteness of the second part of (2.74). We thus arrive at the well-known Kolmogorov–Lamperti tightness result (cf. also [13, Corollary A.11]):
@@ 2.12. Compatible measures in $P_{p}(\mathcal{X})$
The compatibility property was implicitly defined in [9, Section 4] as an admissible property for a collection of transport maps and was further explored in [23]. What we outlined above can be generalized to more than 3 measures and to general spaces, for which we give the following definition, mentioned in [23, Remark 2.3.2]. Now, let $(\mathcal{X},d)$ be a complete separable metric space and $p\in[1,\infty)$.
@@
A prominent and well-known example of compatible measures is Wasserstein geodesics, as highlighted in the remark below. A compatible collection, however, need not lie on a Wasserstein geodesic. Take for instance an arbitrary measure and translate it along a non-geodesic curve. Another example is any collection of measures in $P_{p}(\mathbb{R})$ is compatible. As shown in [23, page 49], any collection of Gaussian measures on $\mathbb{R}^{\mathrm{d}}$ that includes the standard Gaussian distribution and such that the covariance matrices of the measures are simultaneously diagonalizable (i.e. one can diagonalize all with one orthogonal matrix) is compatible. See more nontrivial examples presented in [9, Proposition 4.1] and explained in [23, Section 2.3.2].
@@
If |ji|=1|j-i|=1 or 0, nothing is needed to prove. Let's consider ji>1j-i>1. By triangle inequality and Minkowski's inequality for $L^{p}$ functions, we obtain
@@ 2.13. Further remarks
with a measurable geodesic selection map $\ell:\mathcal{X}\times\mathcal{X}\to Geo([0,1];\mathcal{X})$. We briefly emphasize that such a map exists and that the above is well-defined (using a similar argument as in the proof of [17, Proposition 1]). Define the multi-valued map $L:\mathcal{X}\times\mathcal{X}\to 2^{Geo([0,1];\mathcal{X})}$ as follows:
@@
As $\mathcal{G}(L)$ is closed, it is Borel-measurable, i.e., it belongs to $\mathcal{B}(\mathcal{X}^{2}\times C)$, which is equal to $\mathcal{B}(\mathcal{X}^{2})\otimes\mathcal{C}$, since all spaces are separable here. Obviously, $\mathcal{G}(L)$ is also in $\mathcal{B}(\mathcal{X}^{2})_{\Upsilon}\otimes\mathcal{C}$, where $\mathcal{B}(\mathcal{X}^{2})_{\Upsilon}$ denotes the $\Upsilon$-completion of $\mathcal{B}(\mathcal{X}^{2})$. It follows from a measurable selection theorem by Aumann (see, e.g., [11, Theorem III.22] and [5, page 455]) that there exists a measurable map $\ell:\big(\mathcal{X}^{2},\mathcal{B}(\mathcal{X}^{2})_{\Upsilon}\big)\to\big(C([0,1];\mathcal{X}),\mathcal{C}\big)$ such that $\ell(x,y)\in Geo([0,1];\mathcal{X})$ for $\Upsilon$-a.e. $(x,y)\in\mathcal{X}^{2}$. The above argument for selecting geodesics between two points can easily be generalized to multiple points.
@@ 3.1. Existence of a minimizer for the variational problem
Since the sublevels of $\Psi$ are compact (which follows from lower semi-continuity and relative compactness), the condition (3.2) implies that the family of measures $\{\pi_{n}\}_{n}\subset P(\Gamma_{T})$ is tight by Lemma 2.36. By Prokhorov's theorem, there exists a subsequence $\{\pi_{n_{k}}\}_{k\in\mathbb{N}}$ such that $\pi_{n_{k}}\to\pi$ narrowly on $P(\Gamma_{T})$ as $k\to\infty$. By Lemma 2.43, we have $\pi\in\mathrm{Lift}(\mu_{t})$ and thus $\int_{\Gamma_{T}}\Psi{\mathrm{\,d}}\pi\geq M$. On the other hand, since $\Psi$ is lower semi-continuous, we obtain the reverse inequality
@@ 3.2. From path measures to Wasserstein curves
Next, we show that $t\mapsto\mu_{t}$ is a continuous curve in $p$-Wasserstein space. Fix an arbitrary $t\in[0,T]$ and take a sequence $t_{n}\to t$ as $n\to\infty$. This sequence will eventually fall in $\delta$-neighbourhood of $t$. Thus, by Lebesgue's dominated convergence theorem on the measure space $(\Gamma_{T},\mathcal{B}(\Gamma_{T}),\pi)$ and the integrability assumption (3.5), we obtain
@@
where we used Tonelli's theorem to interchange integrals. Similarly, we have (when $T=1$) that
@@
where we used Beppo Levi's lemma to interchange limit and integral. ∎
@@
by Lebesgue's dominated convergence theorem, where the dominated function
@@ 3.3. From Wasserstein curves to path measures: a superposition principle
$\ignorespaces\ignorespaces(\text{Pr}^{i,i+\frac{2^{n}}{2^{m}}})_{\#}\Upsilon_{n}\in\text{OptCpl}\big(\mu_{t^{(n)}_{i}},\mu_{t^{(n)}_{i+\frac{2^{n}}{2^{m}}}}\big)$,
@@
$=\int_{\bm{\mathcal{X}}_{n}}\Big(\sum_{m=0}^{n}2^{m(\alpha p-1)}\hskip-20.0pt\sum_{\begin{subarray}{c}i=k\frac{2^{n}}{2^{m}}\\ k\in\{0,1,\cdots,2^{m}-1\}\end{subarray}}\hskip-20.0ptd(x_{i},x_{i+\frac{2^{n}}{2^{m}}})^{p}+\frac{2^{n(\alpha p-1)}}{2^{(p-\alpha p)}-1}\sum_{i=0}^{2^{n}-1}d(x_{i},x_{i+1})^{p}\Big){\mathrm{\,d}}\Upsilon_{n}$ (3.60)
$=\sum_{m=0}^{n}2^{m(\alpha p-1)}\hskip-20.0pt\sum_{\begin{subarray}{c}i=k\frac{2^{n}}{2^{m}}\\ k\in\{0,1,\cdots,2^{m}-1\}\end{subarray}}\hskip-20.0pt\int d(x_{i},x_{i+\frac{2^{n}}{2^{m}}})^{p}{\mathrm{\,d}}\Upsilon_{n}+\frac{2^{n(\alpha p-1)}}{2^{(p-\alpha p)}-1}\sum_{i=0}^{2^{n}-1}\int d(x_{i},x_{i+1})^{p}{\mathrm{\,d}}\Upsilon_{n}$ (3.61)
$=\sum_{m=0}^{n}2^{m(\alpha p-1)}\hskip-20.0pt\sum_{\begin{subarray}{c}i=k\frac{2^{n}}{2^{m}}\\ k\in\{0,1,\cdots,2^{m}-1\}\end{subarray}}\hskip-20.0ptW_{p}^{p}(\mu_{t^{(n)}_{i}},\mu_{t^{(n)}_{i+2^{n}/2^{m}}})+\frac{2^{n(\alpha p-1)}}{2^{(p-\alpha p)}-1}\sum_{i=0}^{2^{n}-1}W_{p}^{p}(\mu_{t^{(n)}_{i}},\mu_{t^{(n)}_{i+1}})$ (3.62)
@@
Let us clarify the computation steps. After applying the push-forward (3.56) in the first step (3.59), we note that for fixed $n$ and $\bm{x}=(x_{0},\cdots,x_{N_{n}})$, the curve $t\mapsto\ell_{\bm{x}}(t)$ is piecewise geodesic. We have previously computed the $b^{\alpha,p}$-semi-norm of such curves in (2.55), which we now use to obtain (3.60). The compatibility property (LABEL:eq:2D_marginal_Upsilon) is used to go from (3.61) to (3.62), where even the 2-D marginals of $\Upsilon_{n}$ that are separated by more than one step are given by the Wasserstein distance. The inequality in (3.64) is justified by extending the upper limit of the first sum in (3.63) to infinity, which then corresponds exactly to the definition $|\mu|^{p}_{b^{\alpha,p}}$. The second sum is nothing but one summand in the definition of $|\mu|^{p}_{b^{\alpha,p}}$, so it can be bounded accordingly. To bound (3.65), we use the equivalence (2.23) of the semi-norms $|\cdot|_{W^{\alpha,p}}$ and $|\cdot|_{b^{\alpha,p}}$. The last expression in (3.66) is indeed finite by assumption.
@@
$\int_{\Gamma}\phi(\gamma_{t}){\mathrm{\,d}}\pi(\gamma)\,$ $\overset{(a)}{=}\lim_{k\to\infty}\int_{\Gamma}\phi(\gamma_{[2^{n_{k}}t]/2^{n_{k}}}){\mathrm{\,d}}\pi_{n_{k}}(\gamma)$ (3.70)
$\overset{(b)}{=}\lim_{k\to\infty}\int_{\bm{\mathcal{X}}_{n_{k}}}\phi(x_{[2^{n_{k}}t]}){\mathrm{\,d}}\Upsilon_{n_{k}}(\bm{x})$ (3.71)
$\overset{(c)}{=}\lim_{k\to\infty}\int_{\mathcal{X}}\phi(x){\mathrm{\,d}}\mu_{[2^{n_{k}}t]/2^{n_{k}}}(x)$ (3.72)
$\overset{(d)}{=}\int_{\mathcal{X}}\phi(x){\mathrm{\,d}}\mu_{t}(x).$ (3.73)
@@
where $d(\gamma_{t},\gamma_{[2^{n_{k}}t]/2^{n_{k}}})$ is estimated using Garsia–Rodemich–Rumsey inequality as described in (2.29) and we used Jensen's inequality for the integral. Note that thanks to the calculations in the proof of tightness, we now know that $\int|\gamma|^{p}_{W^{\alpha,p}}{\mathrm{\,d}}\pi_{n_{k}}$ is uniformly finite, which means that $\pi_{n_{k}}$-a.e. $\gamma$ has finite $W^{\alpha,p}$-semi-norm. In the limit, the last expression above approaches zero since
@@
Steps (b)(b)-$(c)$ simply follow from the construction. More precisely, for fixed $k$, the 1-D marginals of $\pi_{n_{k}}$ at time $t$ coincide with $\mu_{t}$ whenever $t$ is of the form $t=\frac{i}{2^{n_{k}}}$ for some integer $i$.
@@
$\overset{(c)}{=}\liminf_{m\to\infty}W_{p}^{p}\big(\mu_{[2^{m}t]/2^{m}},\mu_{[2^{m}s]/2^{m}}\big)$ (3.80)
$\overset{(d)}{=}W_{p}^{p}(\mu_{t},\mu_{s}),$ (3.81)
where in step $(a)$, we applied Fatou's lemma on the measure space $(\Gamma,\mathcal{B}(\Gamma),\pi)$. Notice that for every continuous curve $\gamma\in\Gamma$, we have
@@
where we used Tonelli's theorem. Similarly, we have
@@
where we used Beppo Levi's lemma. This completes the proof. ∎